# Veli: Unsupervised Method and Unified Benchmark for Low-Cost Air Quality Sensor Correction

### v1

@@ Abstract
Urban air pollution is a major health crisis causing millions of premature deaths annually, underscoring the urgent need for accurate and scalable monitoring of air quality (AQ). While low-cost sensors (LCS) offer a scalable alternative to expensive reference-grade stations, their readings are affected by drift, calibration errors, and environmental interference. To address these challenges, we introduce Veli (Reference-free Variational Estimation via Latent Inference), an unsupervised Bayesian model that leverages variational inference to correct LCS readings without requiring co-location with reference stations, eliminating a major deployment barrier. Specifically, Veli constructs a disentangled representation of the LCS readings, effectively separating the true pollutant reading from the sensor noise. To build our model and address the lack of standardized benchmarks in AQ monitoring, we also introduce the Air Quality Sensor Data Repository (AQ-SDR). AQ-SDR is the largest AQ sensor benchmark to date, with readings from 23,737 LCS and reference stations across multiple regions. Veli demonstrates strong generalization across both in-distribution and out-of-distribution settings, effectively handling sensor drift and erratic sensor behavior. Code for model and dataset will be made public when this paper is published.
@@ 1 Introduction
The World Health Organization (WHO) estimated that over 90% of the world’s population breathes air that contains pollutants above WHO guideline levels (World Health Organization 2018). These pollutants are known to cause respiratory and cardiovascular diseases, and are present in high concentrations in urban areas (Zhang et al. 2024). To meet WHO air quality standards, real-time air quality (AQ) monitoring is crucial.
@@
To use LCS to increase the spatial coverage of AQ monitoring, reliable methods for correcting their erratic readings are necessary. Many pre-deployment calibration methods exist for LCS (Delaine, Lebental, and Rivano 2019; Hagan et al. 2018; Maag et al. 2016). However, dense deployment of LCS would require recurrent manual recalibration to prevent issues like sensor drift. To eliminate the need for manual recalibration, numerous works have explored numerical approaches for post-deployment data correction [^1] To avoid confusion, we use the term ‘correction’ for all numerical/algorithmic approaches to data processing, and distinguish it from instrument calibration of the devices.. LCS data correction methods often rely on high-cost reference stations as the ground truth to train supervised machine learning models. A fundamental limitation of these models is their reliance on the co-location of LCS with high-cost stations to collect synchronized data pairs for training, which undermines the core objective of using LCS as an affordable option to increase spatial coverage (Maag, Zhou, and Thiele 2018). Moreover, these data correction models are typically trained over a short period of time (often a few months), making them unreliable for long-term applications due to sensor drift and seasonal variations. Another significant but largely overlooked limitation is that these models often fail to account for real-world operational issues. For instance, deployed LCS exhibit significant bias and drift, and can experience periods of data or connectivity loss, causing their uncorrected readings to mislead end-users and public health analysts (Concas et al. 2021a). Lastly, previous studies do not use a standard benchmark or dataset for model evaluation. The lack of a common benchmark hinders reliable evaluation, as reported metrics often lack the context to compare different methods effectively.
@@ 2.1 Reference-based Methods
A major limitation of pre-deployment reference-based correction methods is the need to co-locate target LCS units next to a reference station for an extended period to collect calibration data, making the deployment of large LCS networks impractical. Moreover, shorter co-location intervals yield models that poorly capture temporal variations such as seasonal changes. Lastly, this initial calibration does not account for long-term sensor drift, necessitating periodic recalibration. The logistical challenges of recalibrating deployed sensors mean that long-term drift often goes uncorrected in many devices. Most early studies adopted simple linear models in the pre-deployment context (e.g., ordinary least squares regression). We refer readers to (Concas et al. 2021a; Maag, Zhou, and Thiele 2018) for a comprehensive review of these methods.
The complexity of LCS errors has recently led to increased interest in non-linear post-deployment correction methods. These methods address the limitations of the traditional design paradigm, which relies on synchronized and co-located LCS-reference pairs, similar to (Ahn et al. 2025). For instance, (Cheng et al. 2019) addressed post-deployment correction using unsynchronized calibration transfer, a technique for in-field calibration via co-location with a reference station. This co-located LCS then serves as an anchor point, providing ground truth for other sensors with no co-located references in the network. While this method reports promising results, it was tested on only seven LCS during a ten-month period. Moreover, the LCS units were deployed in controlled settings, avoiding real-world issues such as missing data and extreme fluctuations. (Wang et al. 2023) proposed CaliFormer, a hybrid reference-based approach combining unsupervised reconstruction with supervised fine-tuning. The model is initially trained to reconstruct the LCS data in an unsupervised manner, and then fine-tuned to correct the results using ground-truth data from the reference stations.
In addition to direct correction of readings, some works have used historical data from reference stations as prior knowledge for LCS correction. Both ‘RHC’ (Li et al. 2020) and the Maximal Correlation Model (Li et al. 2021) leverage historical reference data to align LCS and reference readings’ distributions. A significant limitation is that both approaches were evaluated on short time frames, restricting their applicability for long-term deployments.
@@
A different line of work bypasses LCS correction altogether, creating high-resolution AQ maps by interpolating data directly from a network of reference stations. Both MapTransfer (Cheng et al. 2020) and AirRadar (Wang et al. 2025) interpolate readings from high-cost stations to generate denser pollution maps. Despite their ability to produce high-resolution AQ maps, these approaches depend on reference stations with a sparse deployment across a region, which limits their ability to capture microclimate variations.
@@ 2.2 AQ Benchmarks
In Table 1, we compare previously published datasets and benchmarks that contain LCS data with our new dataset, AQ-SDR. We provide further details on AQ-SDR in Section 4.1 and Appendix C. Previous datasets are either limited to small-scale studies on a regional level (Diez et al. 2024), or cover shorter time periods (Jiao et al. 2016). While some benchmarks provide aligned LCS and reference station readings (Bi et al. 2022; Van Poppel et al. 2023), they do not provide a scale large enough to develop models that can generalize across diverse pollution levels. Our dataset is designed to serve as a unifying benchmark for LCS modeling and correction methods, capturing a wide range of failure modes, distribution shifts, sensor drift, and pollution levels to reflect real-world LCS behavior. AQ-SDR is the largest AQ sensor dataset to date, containing data from 23,737 low-cost and reference sensors across multiple global regions, collected over more than six years of deployment.
@@
(Jiao et al. 2016) 20 10 LCS Only
(Diez et al. 2024) 49 34 LCS Only
(Van Poppel et al. 2023) 85 12 Both
(Bi et al. 2022) 109 22 Both
@@ 3.1 Problem Formulation
A key challenge for reference-free correction is achieving robustness against the diverse failures and environmental factors seen in real-world deployments. Thus, it is essential to use a dataset that contains numerous instances of systematic drift, failures, and other erratic behaviors known to hinder LCS correction when developing and validating correction models (Concas et al. 2021b).
These combined real-world challenges often cause standard denoising and sensor fusion approaches, such as least-squares methods and Kalman filters (Kalman 1960), to fail as their state estimation becomes unreliable when readings are extremely erratic or contain missing values. To further illustrate the issue of sensor drift, we show in Figure 2 an example taken from AQ-SDR, shown as a comparison over a three-year period between a low-cost air quality sensor and a co-located, calibrated reference station in the city of Groningen in the Netherlands. LCS exhibit a noticeable shift in the distribution of their PM2.5 readings over the years, despite having a distribution similar to that of the reference station around its initial deployment in 2019. In contrast, the well-maintained reference station shows consistent behavior, with a nearly identical data distribution over the same period.
@@
To accurately model sensor bias, we build upon established findings that show LCS errors exhibit both nonlinear patterns and a systematic bias with heteroscedastic variance caused by environmental factors (Sharma et al. 2025; Concas et al. 2021b). These error characteristics motivated the use of techniques like Gaussian process regression to correct LCS readings (Malings et al. 2019; Li et al. 2023). Consequently, we adopt a similar rationale and propose a probabilistic sensor fusion model based on advances in Variational Autoencoders (VAEs) (Kingma and Welling 2014).
@@ 3.2 Model Overview
Figure 3: Veli structure following the derivation in Section 3.3. The input starts with AQ readings $x$ and auxiliary mask of ‘NA’ readings $\psi$ on the left, propagating through the model’s layers to generate a prediction of clean readings $\hat{y}$. Conditioning on $\psi$ is omitted in some blocks for visual clarity but is implemented properly. Prior distribution blocks (green) are used in the training to estimate the variational distribution blocks (blue), which are used in the inference as indicated by the blue dashed line. All distribution blocks are modeled by two multilayer perceptron (MLP) layers followed by an MLP layer for each of the mean and variance. The losses $\mathcal{L}_{KL_{z}}$, $\mathcal{L}_{KL_{y}}$, and $\mathcal{L}_{recon}$ correspond to the three terms in eq. (6). Sampling refers to the traditional reparameterization in VAEs (Kingma and Welling 2014).
In this implementation, we focus on correcting individual snapshots of LCS readings rather than modeling changes over time. We propose this design decision for two key reasons: First, it is difficult to obtain perfectly time-aligned data streams from multiple adjacent sensors without encountering gaps or simultaneous failures. Second, simultaneously modeling time alongside all noise patterns (e.g., spikes, missing data) compromises the model’s ability to capture diverse non-temporal noise patterns. While our model processes hourly readings per pass, this snapshot-based approach does not discard the underlying temporal information. Since the correction model uses Lipschitz continuous layers (MLPs), temporal signatures in the corrected output remain preserved, as previously explored in (Virmaux and Scaman 2018).
@@ 3.3 LCS Noise Model
In this section, we provide the necessary formulation to build Veli. We refer readers to (Kingma and Welling 2014) for more insights on the foundations of VAEs, and provide a more detailed derivation in Appendix A.
@@
Standard VAEs typically use a standard Gaussian prior, $N(0,I)$. However, to build a more identifiable and flexible prior, we follow the approach in (Khemakhem et al. 2020) and introduce $\psi$ as our auxiliary parameter. This approach allows the latent space to effectively learn diverse variations within the input data, which is essential in filtering erratic behavior. In the same manner, and since we are operating without LCS-reference-paired readings $(x_{\rm noise},y)$, we treat $y$ as a latent variable whose prior distribution is given as:
@@ 3.4 Variational Approximations
To estimate the distributions of $y$ and $z$ through the term $p(y,z\mid\psi)$, we will need to evaluate an intractable integral with no closed-form solution. Therefore, we introduce approximate variational distributions similar to (Kingma and Welling 2014), defined as:
@@
$\log p(x_{\rm noise}\mid\psi)\geq\mathbb{E}$ [${}_{\,q_{\rm\theta,\phi}(y,z\mid x_{\rm noise},\psi)}\bigl{[}\log p(x_{\rm noise},y,z\mid\psi)$
@@
where $\hat{y}$ is the sampled prediction from the distribution in eq. (5) during training, but is taken as a point estimate during inference. Here $\alpha,\beta_{\rm z},\beta_{\rm y}>0$ are tunable coefficients similar to (Higgins et al. 2017).
@@ 4.1 AQ-SDR Dataset
To build our model, we use our proposed dataset, the AQ-SDR, which aggregates the LCS data from three major citizen-science initiatives: SamenMeten, Sensor.Community, and Location Aware Sensing System (LASS) community (Chen et al. 2017). To supply reference measurements to validate our method, we provide data from four authoritative sources: LuchtMeetNet (Air Measurement Network), the Royal Netherlands Meteorological Institute, the European Environment Agency, and the Taiwanese Ministry of Environment Open Data. The majority of the deployed LCS began operating in 2019 and continue to provide measurements, with a smaller subset having been operational since before 2019. To further evaluate our model’s generalizability across different pollution levels and regions, we create two partitions of the AQ-SDR dataset: an in-distribution set and an out-of-distribution set. We illustrate the difference between the in-distribution data and out-of-distribution data using two samples shown in Figure 4. The left-skewed distribution of the in-distribution data (Netherlands) reflects lower pollution levels, in contrast to the right-skewed distribution of the out-of-distribution data (Taiwan), which indicates higher pollution levels.
@@ 4.2 Implementation Details & Evaluation Metrics
All MLP layers that do not concatenate inputs use a hidden dimension of 32. For out-of-distribution fine-tuning, we froze the decoder and trained only the encoder for an additional 30 epochs on the new data distribution. Other implementation, tensor preparation, and evaluation details match previous works in time-series modeling (Liu et al. 2024).
To evaluate our model, we use Mean Absolute Error (MAE) as the standard metric from the literature to compare model outputs to reference readings (Concas et al. 2021a; Maag, Zhou, and Thiele 2018). We also report the average and standard deviation of the output over five runs with different random seeds for all numerical results. In all experiments, ‘raw LCS’ is the input to Veli.
As this is the first work to propose a completely reference-free correction method, no other methods exist for a direct comparison. In addition, no large-scale unifying benchmark exists aside from AQ-SDR. We instead provide a comparison against traditional blind denoising techniques like Kalman Filters (KF) (Kalman 1960) and Principal Component Analysis (PCA) denoising (Weston, Schölkopf, and Bakir 2003). A KNN imputer was used to enable these two methods to run on data with missing readings. In addition, we provide extended results and analysis in Appendix B.

### v2

@@ Abstract
Urban air pollution is a major health crisis causing millions of premature deaths annually, underscoring the urgent need for accurate and scalable monitoring of air quality (AQ). While low-cost sensors (LCS) offer a scalable alternative to expensive reference-grade stations, their readings are affected by drift, calibration errors, and environmental interference. To address these challenges, we introduce Veli (Reference-free Variational Estimation via Latent Inference), an unsupervised Bayesian model that leverages variational inference to correct LCS readings without requiring co-location with reference stations, eliminating a major deployment barrier. Specifically, Veli constructs a disentangled representation of the LCS readings, effectively separating the true pollutant reading from the sensor noise. To build our model and address the lack of standardized benchmarks in AQ monitoring, we also introduce the Air Quality Sensor Data Repository (AQ-SDR). AQ-SDR is the largest AQ sensor benchmark to date, with readings from 23,737 LCS and reference stations across multiple regions. Veli demonstrates strong generalization across both in-distribution and out-of-distribution settings, effectively handling sensor drift and erratic sensor behavior. Code for model and dataset will be made public when this paper is published. The appendices are available in the extended version.
@@ 1 Introduction
The World Health Organization (WHO) estimated that over 90% of the world’s population breathes air that contains pollutants above WHO guideline levels (WHO2025HealthRisks). These pollutants are known to cause respiratory and cardiovascular diseases, and are present in high concentrations in urban areas (Zhang2025PM2_5CVD). To meet WHO air quality standards, real-time air quality (AQ) monitoring is crucial.
@@
To use LCS to increase the spatial coverage of AQ monitoring, reliable methods for correcting their erratic readings are necessary. Many pre-deployment calibration methods exist for LCS (Delaine2019Calibration; Hagan2018Sensors; Maag2016PreDeployment). However, dense deployment of LCS would require recurrent manual recalibration to prevent issues like sensor drift. To eliminate the need for manual recalibration, numerous works have explored numerical approaches for post-deployment data correction [^1] To avoid confusion, we use the term ‘correction’ for all numerical/algorithmic approaches to data processing, and distinguish it from instrument calibration of the devices.. LCS data correction methods often rely on high-cost reference stations as the ground truth to train supervised machine learning models. A fundamental limitation of these models is their reliance on the co-location of LCS with high-cost stations to collect synchronized data pairs for training, which undermines the core objective of using LCS as an affordable option to increase spatial coverage (survey2). Moreover, these data correction models are typically trained over a short period of time (often a few months), making them unreliable for long-term applications due to sensor drift and seasonal variations. Another significant but largely overlooked limitation is that these models often fail to account for real-world operational issues. For instance, deployed LCS exhibit significant bias and drift, and can experience periods of data or connectivity loss, causing their uncorrected readings to mislead end-users and public health analysts (Concas2021LowCost). Lastly, previous studies do not use a standard benchmark or dataset for model evaluation. The lack of a common benchmark hinders reliable evaluation, as reported metrics often lack the context to compare different methods effectively.
@@ 2.1 Reference-based Methods
A major limitation of pre-deployment reference-based correction methods is the need to co-locate target LCS units next to a reference station for an extended period to collect calibration data, making the deployment of large LCS networks impractical. Moreover, shorter co-location intervals yield models that poorly capture temporal variations such as seasonal changes. Lastly, this initial calibration does not account for long-term sensor drift, necessitating periodic recalibration. The logistical challenges of recalibrating deployed sensors mean that long-term drift often goes uncorrected in many devices. Most early studies adopted simple linear models in the pre-deployment context (e.g., ordinary least squares regression). We refer readers to (Concas2021LowCost; survey2) for a comprehensive review of these methods.
The complexity of LCS errors has recently led to increased interest in non-linear post-deployment correction methods. These methods address the limitations of the traditional design paradigm, which relies on synchronized and co-located LCS-reference pairs, similar to (tesla). For instance, (Cheng2019ICT) addressed post-deployment correction using unsynchronized calibration transfer, a technique for in-field calibration via co-location with a reference station. This co-located LCS then serves as an anchor point, providing ground truth for other sensors with no co-located references in the network. While this method reports promising results, it was tested on only seven LCS during a ten-month period. Moreover, the LCS units were deployed in controlled settings, avoiding real-world issues such as missing data and extreme fluctuations. (Wang2023CaliFormer) proposed CaliFormer, a hybrid reference-based approach combining unsupervised reconstruction with supervised fine-tuning. The model is initially trained to reconstruct the LCS data in an unsupervised manner, and then fine-tuned to correct the results using ground-truth data from the reference stations.
In addition to direct correction of readings, some works have used historical data from reference stations as prior knowledge for LCS correction. Both ‘RHC’ (rch2020) and the Maximal Correlation Model (maximizingcorrelation) leverage historical reference data to align LCS and reference readings’ distributions. A significant limitation is that both approaches were evaluated on short time frames, restricting their applicability for long-term deployments.
@@
A different line of work bypasses LCS correction altogether, creating high-resolution AQ maps by interpolating data directly from a network of reference stations. Both MapTransfer (Cheng2020MapTransfer) and AirRadar (Wang2025AirRadar) interpolate readings from high-cost stations to generate denser pollution maps. Despite their ability to produce high-resolution AQ maps, these approaches depend on reference stations with a sparse deployment across a region, which limits their ability to capture microclimate variations.
@@ 2.2 AQ Benchmarks
In Table 1, we compare previously published datasets and benchmarks that contain LCS data with our new dataset, AQ-SDR. We provide further details on AQ-SDR in Section 4.1 and Appendix C. Previous datasets are either limited to small-scale studies on a regional level (Diez2024QUANT), or cover shorter time periods (Jiao2016CAIRSENSE). While some benchmarks provide aligned LCS and reference station readings (Bi2022; VanPoppel2023SensEURCity), they do not provide a scale large enough to develop models that can generalize across diverse pollution levels. Our dataset is designed to serve as a unifying benchmark for LCS modeling and correction methods, capturing a wide range of failure modes, distribution shifts, sensor drift, and pollution levels to reflect real-world LCS behavior. AQ-SDR is the largest AQ sensor dataset to date, containing data from 23,737 low-cost and reference sensors across multiple global regions, collected over more than six years of deployment.
@@
(Jiao2016CAIRSENSE) 20 10 LCS Only
(Diez2024QUANT) 49 34 LCS Only
(VanPoppel2023SensEURCity) 85 12 Both
(Bi2022) 109 22 Both
@@ 3.1 Problem Formulation
A key challenge for reference-free correction is achieving robustness against the diverse failures and environmental factors seen in real-world deployments. Thus, it is essential to use a dataset that contains numerous instances of systematic drift, failures, and other erratic behaviors known to hinder LCS correction when developing and validating correction models (lcs_outdoor_survey).
These combined real-world challenges often cause standard denoising and sensor fusion approaches, such as least-squares methods and Kalman filters (kalman), to fail as their state estimation becomes unreliable when readings are extremely erratic or contain missing values. To further illustrate the issue of sensor drift, we show in Figure 2 an example taken from AQ-SDR, shown as a comparison over a three-year period between a low-cost air quality sensor and a co-located, calibrated reference station in the city of Groningen in the Netherlands. LCS exhibit a noticeable shift in the distribution of their PM2.5 readings over the years, despite having a distribution similar to that of the reference station around its initial deployment in 2019. In contrast, the well-maintained reference station shows consistent behavior, with a nearly identical data distribution over the same period.
@@
To accurately model sensor bias, we build upon established findings that show LCS errors exhibit both nonlinear patterns and a systematic bias with heteroscedastic variance caused by environmental factors (lcs_variance; lcs_outdoor_survey). These error characteristics motivated the use of techniques like Gaussian process regression to correct LCS readings (lcs_gaussian; variationalbias). Consequently, we adopt a similar rationale and propose a probabilistic sensor fusion model based on advances in Variational Autoencoders (VAEs) (vae).
@@ 3.2 Model Overview
Figure 3: Veli structure following the derivation in Section $3.$ The input starts with AQ readings $x$ and auxiliary mask of ‘NA’ readings $\psi$ on the left, propagating through the model’s layers to generate a prediction of clean readings $\hat{y}$. Conditioning on $\psi$ is omitted in some blocks for visual clarity but is implemented properly. Prior distribution blocks (green) are used in the training to estimate the variational distribution blocks (blue), which are used in the inference as indicated by the blue dashed line. All distribution blocks are modeled by two multilayer perceptron (MLP) layers followed by an MLP layer for each of the mean and variance. The losses $\mathcal{L}_{KL_{z}}$, $\mathcal{L}_{KL_{y}}$, and $\mathcal{L}_{recon}$ correspond to the three terms in eq. (6). Sampling refers to the traditional reparameterization in VAEs (vae).
In this implementation, we focus on correcting individual snapshots of LCS readings rather than modeling changes over time. We propose this design decision for two key reasons: First, it is difficult to obtain perfectly time-aligned data streams from multiple adjacent sensors without encountering gaps or simultaneous failures. Second, simultaneously modeling time alongside all noise patterns (e.g., spikes, missing data) compromises the model’s ability to capture diverse non-temporal noise patterns. While our model processes hourly readings per pass, this snapshot-based approach does not discard the underlying temporal information. Since the correction model uses Lipschitz continuous layers (MLPs), temporal signatures in the corrected output remain preserved, as previously explored in (lipschitz).
@@ 3.3 LCS Noise Model
In this section, we provide the necessary formulation to build Veli. We refer readers to (vae) for more insights on the foundations of VAEs, and provide a more detailed derivation in Appendix A.
@@
Standard VAEs typically use a standard Gaussian prior, $N(0,I)$. However, to build a more identifiable and flexible prior, we follow the approach in (ivae) and introduce $\psi$ as our auxiliary parameter. This approach allows the latent space to effectively learn diverse variations within the input data, which is essential in filtering erratic behavior. In the same manner, and since we are operating without LCS-reference-paired readings $(x_{\rm noise},y)$, we treat $y$ as a latent variable whose prior distribution is given as:
@@ 3.4 Variational Approximations
To estimate the distributions of $y$ and $z$ through the term $p(y,z\mid\psi)$, we will need to evaluate an intractable integral with no closed-form solution. Therefore, we introduce approximate variational distributions similar to (vae), defined as:
@@
$\log p(x_{\rm noise}\mid\psi)\geq\mathbb{E}$ $_{\,q_{\rm\theta,\phi}(y,z\mid x_{\rm noise},\psi)}\bigl[\log p(x_{\rm noise},y,z\mid\psi)$
@@
where $\sigma_{\mathrm{sens}}^{2}(z)_{i}$ and $\mu_{\mathrm{sens}}(z)_{i}$ are the non-linear bias and heteroscedastic terms in eq. (1), and $\hat{y}$ is the sampled prediction from the distribution in eq. (5) during training, but is taken as a point estimate during inference. Here $\alpha,\beta_{\rm z},\beta_{\rm y}>0$ are tunable coefficients similar to (betavae).
@@ 4.1 AQ-SDR Dataset
To build our model, we use our proposed dataset, the AQ-SDR, which aggregates the LCS data from three major citizen-science initiatives: SamenMeten, Sensor.Community, and Location Aware Sensing System (LASS) community (taiwan_data). To supply reference measurements to validate our method, we provide data from four authoritative sources: LuchtMeetNet (Air Measurement Network), the Royal Netherlands Meteorological Institute, the European Environment Agency, and the Taiwanese Ministry of Environment Open Data. The majority of the deployed LCS began operating in 2019 and continue to provide measurements, with a smaller subset having been operational since before 2019. To further evaluate our model’s generalizability across different pollution levels and regions, we create two partitions of the AQ-SDR dataset: an in-distribution set and an out-of-distribution set. We illustrate the difference between the in-distribution data and out-of-distribution data using two samples shown in Figure 4. The left-skewed distribution of the in-distribution data (Netherlands) reflects lower pollution levels, in contrast to the right-skewed distribution of the out-of-distribution data (Taiwan), which indicates higher pollution levels.
@@ 4.2 Implementation Details & Evaluation Metrics
All MLP layers that do not concatenate inputs use a hidden dimension of 32. For out-of-distribution fine-tuning, we froze the decoder and trained only the encoder for an additional 30 epochs on the new data distribution. Other implementation, tensor preparation, and evaluation details match previous works in time-series modeling (liu2024itransformer).
To evaluate our model, we use Mean Absolute Error (MAE) as the standard metric from the literature to compare model outputs to reference readings (Concas2021LowCost; survey2). We also report the average and standard deviation of the output over five runs with different random seeds for all numerical results. In all experiments, ‘raw LCS’ is the input to Veli.
As this is the first work to propose a completely reference-free correction method, no other methods exist for a direct comparison. In addition, no large-scale unifying benchmark exists aside from AQ-SDR. We instead provide a comparison against traditional blind denoising techniques like Kalman Filters (KF) (kalman) and Principal Component Analysis (PCA) denoising (pca). A KNN imputer was used to enable these two methods to run on data with missing readings. In addition, we provide extended results and analysis in Appendix B.